# Automaton Constrained Q-Learning

### v1

@@ 1 Introduction
Despite the success of Goal-Conditioned Reinforcement Learning (GCRL) [1, 2] and Safe RL [3, 4], these approaches fall short in addressing the combined challenges of sequential goals and dynamic safety constraints. GCRL is effective for reaching individual goals but lacks mechanisms for reasoning over goal sequences or ensuring safety. Safe RL enforces fixed safety constraints during training, but typically assumes they remain static throughout execution, limiting its applicability to temporally evolving tasks.
@@
Formal languages such as LTL [5] provide an expressive framework for specifying complex tasks involving multiple interdependent goals and non-stationary safety constraints. However, reward functions defined directly from LTL specifications are inherently non-Markovian, since temporal properties refer to entire trajectories rather than individual transitions. This violates the Markov assumption—a fundamental premise of standard RL—which assumes that transitions and rewards depend only on the current state and action. As a result, RL algorithms built on the Markov Decision Process (MDP) formalism [6, 7] are ill-suited for direct application to LTL-defined objectives.
@@
To overcome this, a few studies [8, 9, 10] have proposed to translate the LTL formula into an automaton that tracks temporal progress toward task completion and enables defining Markovian rewards corresponding to temporal specifications. While this approach allows for potentially handling arbitrary LTL formulas, its generality is tied to the use of sparse binary rewards derived from the Boolean evaluation of the specification. In complex environments, however, such sparse signals are encountered too infrequently to effectively guide behavior, and designing denser rewards typically requires additional domain knowledge not readily available in many robotics tasks. Furthermore, these methods only implicitly deal with ensuring safety, as safety properties are fundamental components of many LTL specifications [11], and typically rely on ad-hoc mechanisms, such as halting rollouts [12] and reward shaping [13], both of which tend to be brittle and ineffective in complex domains. Other approaches seek to improve scalability by employing hierarchical [13] or goal-conditioned [14] policies but sacrifice generality, particularly with respect to safety constraints.
@@
To bridge the gap between LTL-capable algorithms that scale poorly and scalable RL algorithms that lack support for most LTL tasks, we propose ACQL, which lifts Safe RL and GCRL to the class of problems expressible as LTL formulae in the recurrence class [15]. This algorithm, which we consider our primary contribution, is built on top of two technical novelties that address distinct challenges associated with LTL problems. First, to overcome the poor scalability of sparse rewards, we encode automaton states with their associated goals, enabling the use of goal-conditioned techniques such as Hindsight Experience Replay (HER) [2] to densify reward signals. Second, inspired by Hamilton-Jacobi (HJ) reachability analysis [16, 17], we employ a minimum-safety-based product Constrained Markov Decision Process (CMDP) formulation that enforces compliance with arbitrary LTL safety constraints at optimality. We demonstrate these technical contributions are necessary in enabling our ACQL algorithm to significantly outperform other algorithms for solving LTL tasks. Additionally, we demonstrate its effectiveness in learning real-world-deployable policies for a 6-DOF robot arm operating in a storage cabinet environment.
@@ 2 Related Work
The use of LTL specifications with RL algorithms has been a popular topic in recent years [18, 19, 10, 13, 20, 21]. Some works directly define non-Markovian reward functions based on overall task satisfaction, but these methods can only support LTL formulae that are satisfiable over finite prefixes [18, 22, 23, 24]. In restricted cases where task satisfaction is differentiable with respect to a policy’s actions, it is also possible to directly train sequence models that satisfy the specification [21]. The most popular and general approaches, however, convert LTL tasks into automata to use in conjunction with standard deep RL techniques for MDP [8, 25, 26, 19, 27, 28, 29, 9, 10]. We specifically aim to improve on the scalability of these methods for solving tasks in high-dimensional environments. Of these techniques, we specifically highlight Reward Machines [30] as a framework with similar generality to our method and the first baseline we compare against.
@@
Our work utilizes techniques from GCRL [31, 32] to accelerate learning for LTL tasks. Other methods have taken a related approach by learning or re-using goal-conditioned policies that are hierarchically guided by an LTL expression to enable zero-shot generalization to new tasks [14]. Using discrete skills as opposed to a goal-conditioned policy has also been explored in [20, 13]. All these approaches require the user to train a different skill for every automaton edge or a different goal-conditioned policy for every safety constraint that appears in the task, except for limited classes of safety constraints. This is impractical for most real-world tasks and, in particular, when safety constraints are changing throughout a task. Conversely, our method is able to learn a single goal-conditioned policy for all goals in a particular LTL expression while still accounting for arbitrary, potentially non-stationary safety constraints. While ACQL has been developed for online RL rather than zero-shot generalization, there is a significant overlap in the tasks that both can ultimately solve. Therefore, we include the Logical Options Framework (LOF) [13] as a representative for these methods in our experiments.
@@
Our work is also focused on solving LTL tasks that feature safety constraints. Safety constraints are often indirectly addressed in the methods discussed above. They primarily focus on making unsafe actions suboptimal by either terminating rollouts [12, 8, 30, 29] or applying negative reward shaping [13, 25] when an unsafe automaton transition occurs. Other methods seek to handle safety more robustly by explicitly requiring discrete state spaces or system dynamics knowledge to ensure safe transitions [9, 33]. In contrast, techniques from the Safe RL literature can robustly enforce stationary safety constraints in complex environments without such assumptions [3, 34, 35, 36, 37, 4, 38]. Our approach builds on these Safe RL methods to address non-stationary safety constraints induced by LTL specifications.
@@ 3 Preliminaries
RL problems with constraints are typically modeled as discounted CMDP, which are defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},d_{0},r,c,L,\gamma)$. This tuple consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow P(\mathcal{S})$ (where $P(\mathcal{S})$ represents the space of probability measures over $\mathcal{S}$), an initial state distribution $d_{0}\in P(\mathcal{S})$, reward and constraint functions $r,\;c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, a discount factor $\gamma$, and a limit for constraint violation $L\in\mathbb{R}$ [3]. Given a CMDP, the standard objective is to find the stationary policy $\pi:\mathcal{S}\to P(\mathcal{A})$ satisfying the constrained maximization
$\max_{\pi}\;J_{r}(\pi)\quad\textnormal{s.t.}\quad J_{c}(\pi)<L,$ (1)
@@
LTL [5] is an extension of propositional logic for reasoning about systems over time. LTL formulae $\phi\in\Phi$ consist of atomic propositions $p$ from a set $AP$, the standard boolean operators “not” ($\neg$), “and” ($\wedge$), and “or” ($\vee$), and temporal operators that reference the value of propositions in the future. These operators are “next” ($\circ$), “eventually” ($\lozenge$), and “always” ($\square$). We follow the definitions for these operators given in [39]. In our algorithm, we use Signal Temporal Logic (STL), which further extends LTL with quantitative semantics and requires that every atomic proposition $p\in AP$ is defined as a real-valued function of the system state. The quantitative semantics is defined by a function $\rho:\Sigma^{\omega}\times\Phi\rightarrow\mathbb{R}$ that produces a robustness value representing how much a signal $\omega\in\Sigma^{\omega}$ satisfies or violates the property specified by an STL formula $\phi\in\Phi$. Our exact implementation of these quantitative semantics is based on [40]. Every temporal logic constraint can be expressed as the conjunction of a “safety” and “liveness” constraint [41], where a safety constraint is informally defined as a requirement that something must never happen for the task to succeed, and a liveness constraint is defined as a requirement that something must happen for the task to succeed. We take advantage of this dichotomy in our method.
@@
The robotics tasks discussed so far can be expressed more specifically as STL formulae in the recurrence class, as defined in [15], and all such formulae can be translated into abstract machines called Deterministic Büchi Automata [39]. A DBA is formally defined by the tuple $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$, consisting of an alphabet $\Sigma$, a set of internal states $\mathcal{Q}$, a transition function $\delta:\mathcal{Q}\times\Sigma\rightarrow\mathcal{Q}$, an initial state $q_{0}\in\mathcal{Q}$, and a set of accepting states $F\subseteq\mathcal{Q}$. An automaton is used to process an infinite sequence of arbitrary symbols from the alphabet $\Sigma$ and determine whether the sequence satisfies or does not satisfy the original logical expression. In our setting, a symbol processed by the automaton is an MDP state $s\in\mathcal{S}$. The automaton $A$ is always in some state $q\in\mathcal{Q}$, and starts in the state $q_{0}$. Each element $\sigma\in\Sigma$ of the input sequence causes some change in the automaton’s internal state according to the transition function $\delta$. For convenience, we will refer to edges in the automaton with transition predicates. A transition predicate is a propositional formula that holds for all $\sigma\in\Sigma$ which induce a transition between two states $q_{i},q_{j}\in\mathcal{Q}$ according to $\delta$. By processing each state while an agent interacts with an MDP, the internal automaton state provides information to the agent about its current progress in the task. This is the fundamental idea behind a Product MDP, which combines any MDP with a deterministic automaton and is a standard technique in model checking for probabilistic systems [39]. A complete definition for a product MDP in an RL context can be found in previous work such as [10, 30].
@@ 4 Method
Figure 1: The ACQL algorithm relies on a novel augmented formulation of CMDP (left). An input task specification $\phi$ is converted into the DBA AA, from which safety constraints and subgoals are collected into mappings $S$ and $G$ respectively. The learning agent receives subgoals $g_{1},\cdots,g_{n}$ at every stage in the task from $G$ and safety constraint feedback in $c^{A}$ from $S$. Trajectories induced by the policy $\pi_{j}$ are collected into a replay buffer $R$, from which batches $\mathcal{B}^{\tau}_{j}$ are sampled and modified using HER [2]. From these modified trajectories, mini-batches $\mathcal{B}_{i}$ of transitions $(s^{A}_{t},a_{t},r^{A}_{t},c^{A}_{t},s^{A}_{t+1})$ are used to compute the targets $y_{t}^{r}$ in (3) and $y_{t}^{c}$ in (4) for training models of the state-action value function and safety function, $Q^{r}_{\theta}$ and $Q^{c}_{\psi}$, from which an updated policy $\pi_{j+1}$ is derived.
@@ 4.1 Augmented Product CMDP Formulation
To construct the augmented product CMDP used by our algorithm, we begin by translating the input STL task specification $\phi$ into a DBA $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$ using the SPOT library [42]. As described in Section 3, this automaton represents the structure of the task. To better guide learning, we extract from this automaton additional structure that captures the task’s safety and liveness requirements. Intuitively, these represent conditions that must always hold to avoid failure (safety) and conditions that must eventually hold to make progress toward task completion (liveness) [41, 11]. This information is summarized in two mappings: a safety condition mapping $S:\mathcal{Q}\rightarrow\Phi$, which assigns to each automaton state a proposition that must remain true, and a liveness condition mapping $O:\mathcal{Q}\rightarrow\Phi$, which assigns to each automaton state a proposition that must be eventually satisfied to proceed in the task. Based on $O$, we define $G:\mathcal{Q}\rightarrow\mathcal{G}^{+}$ to explicitly link automaton states to subgoals, where $\mathcal{G}^{+}=\bigcup_{i=1}^{n}\mathcal{G}^{i}$ represents the full space of possible subgoal lists up to a task-dependent length $n$.
@@
Second, to obtain liveness constraints, we ignore the above transitions for safety conditions and examine the remaining predicates guarding outgoing transitions from each state $q$. These can only specify the conditions necessary to make progress in the task. Therefore, we can take the disjunction of these transition predicates to obtain the formula $O(q)$ that must be satisfied in order to transition beyond $q$. Following prior work on learning LTL tasks [13, 14], we assume that a subset of the atomic propositions, $AP_{\text{subgoal}}\subseteq AP$, represent subgoal propositions and are parameterized by values $g$ within a subgoal space $\mathcal{G}\subseteq\mathcal{S}$. Note that we consider only a subset since not all the propositions in our modified DBA are necessarily related to achieving subgoals. Using this fact, we filter the predicates in $O(q)$ to retain only those associated with these subgoal propositions, which in turn provides a list of subgoals $G(q)=g_{1},\dots,g_{n}$ that are relevant to the logical formula given by $O(q)$. An illustrative example for these two steps is provided in the appendix.
@@
Now, let $\mathcal{M}^{A}=(\mathcal{S}^{A},\mathcal{A}^{A},\mathcal{T}^{A},d_{0}^{A},r^{A},c^{A},\gamma,L)$ be a new CMDP formed by augmenting the original MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},d_{0},r,\gamma)$ with the DBA $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$. The new state space $\mathcal{S}^{A}=\mathcal{S}\times\mathcal{G}^{+}\times\mathcal{Q}$ is the Cartesian product of the original state space $\mathcal{S}$, the space for every possible list of subgoals $\mathcal{G}^{+}$, and the set of automaton states $\mathcal{Q}$. A state $s^{A}\in\mathcal{S}^{A}$ can be written as $\langle s,g^{+},q\rangle$, where $s\in\mathcal{S}$, $g^{+}\in\mathcal{G}^{+}$, and $q\in\mathcal{Q}$ are the constituent MDP state, goal-list and automaton state of $s^{A}$. The action space $\mathcal{A}^{A}=\mathcal{A}$ is unmodified. The new transition dynamics $\mathcal{T}^{A}$ are defined so that a transition to $\langle s^{\prime},g^{+}{{}^{\prime}},q^{\prime}\rangle\in\mathcal{S}^{A}$ from $\langle s,g^{+},q\rangle\in\mathcal{S}^{A}$ is impossible if the automaton does not support a transition from $q$ to $q^{\prime}$ when entering the state $s^{\prime}$; i.e.,
$\mathcal{T}^{A}$ $(\langle s^{\prime},g^{+}{{}^{\prime}},q^{\prime}\rangle|\langle s,g^{+},q\rangle,a)=\begin{cases}\mathcal{T}(s^{\prime}|s,a)&\text{if}\ q^{\prime}=\delta(q,s^{\prime}),g^{+}{{}^{\prime}}=G(q^{\prime}),\\ 0&\text{otherwise}.\end{cases}$

### v2

@@ 1 Introduction
Despite the success of Goal-Conditioned Reinforcement Learning (GCRL) [schaul2015uvfa, andrychowicz2017her] and Safe RL [altman1999constrained, achiam2017cpo], these approaches fall short in addressing the combined challenges of sequential goals and dynamic safety constraints. GCRL is effective for reaching individual goals but lacks mechanisms for reasoning over goal sequences or ensuring safety. Safe RL enforces fixed safety constraints, but typically assumes they remain static throughout the task, limiting its applicability to temporally evolving tasks.
@@
Formal languages such as LTL [pnueli1977ltl] provide an expressive framework for specifying complex tasks involving multiple interdependent goals and non-stationary safety constraints. However, reward functions defined directly from LTL specifications are inherently non-Markovian, since temporal properties refer to entire trajectories rather than individual transitions. This violates the Markov assumption—a fundamental premise of standard RL—which assumes that transitions and rewards depend only on the current state and action. As a result, RL algorithms built on the Markov Decision Process (MDP) formalism [haarnoja2018soft, schulman2017proximalpolicyoptimizationalgorithms] are ill-suited for direct application to LTL-defined objectives.
@@
To overcome this, some studies [camacho2019ltl, hasanbeig2020cautious, hasanbeig2020deep] have proposed to translate the LTL formula into an automaton that tracks temporal progress toward task completion and enables defining Markovian rewards corresponding to temporal specifications. While this approach allows for potentially handling arbitrary LTL formulas, its generality is tied to the use of sparse binary rewards derived from the Boolean evaluation of the specification. In complex environments, however, such sparse signals are encountered too infrequently to effectively guide behavior, and designing denser rewards typically requires additional domain knowledge not readily available in many robotics tasks. Furthermore, these methods only implicitly deal with ensuring safety, as safety properties are fundamental components of many LTL specifications [alpern1987Recognizing], and typically rely on ad-hoc mechanisms, such as halting rollouts [toroicarte2018RewardMachines] and reward shaping [araki2021logicaloptionsframework], both of which tend to be brittle and ineffective in complex domains. Other approaches seek to improve scalability by employing hierarchical [araki2021logicaloptionsframework] or goal-conditioned [qiu2023instructing] policies but sacrifice generality, particularly with respect to safety constraints.
@@
To bridge the gap between LTL-capable algorithms that scale poorly and scalable RL algorithms that lack support for most LTL tasks, we propose ACQL, which lifts Safe RL and GCRL to the class of problems expressible as LTL formulae in the recurrence class [manna1990hierarchy]. This algorithm, which we consider our primary contribution, is built on top of two technical novelties that address distinct challenges associated with LTL problems. First, to overcome the poor scalability of sparse rewards, we encode automaton states with their associated goals, enabling the use of goal-conditioned techniques such as Hindsight Experience Replay (HER) [andrychowicz2017her] to densify reward signals. Second, inspired by Hamilton-Jacobi (HJ) reachability analysis [fisac2019bridging, hsu2021safety], we employ a minimum-safety-based product Constrained Markov Decision Process (CMDP) formulation that enforces compliance with arbitrary LTL safety constraints at optimality. We demonstrate these technical contributions are necessary in enabling our ACQL algorithm to significantly outperform other algorithms for solving LTL tasks. Additionally, we demonstrate its effectiveness in learning real-world-deployable policies for a 6-DOF robot arm operating in a storage cabinet environment.
@@ 2 Related Work
The use of LTL specifications with RL algorithms has been a popular topic in recent years [li2017reinforcement, li2019formal, hasanbeig2020deep, vaezipoor2021ltl2action, araki2021logicaloptionsframework, tasse2022skill, yalcinkaya2023automata, xiong2024dscrl, xu2024generalization, yalcinkaya2024compositional, yalcinkaya2025automataconditionedrl, jackermeier2025deepltl, shah2025ltlconstrained]. One approach to this topic has been directly defining non-Markovian reward functions based on overall task satisfaction, but this only supports LTL formulae that are satisfiable over finite prefixes [li2017reinforcement, li2018policysearch, aksaray2016qlearningrobustsatisfactionsignal, balakrishnan2019structured]. In restricted cases where task satisfaction is differentiable with respect to a policy’s actions, it is also possible to directly train sequence models that satisfy the specification [xiong2024dscrl]. The most popular and general methods, however, convert LTL tasks into automata to use in conjunction with standard deep RL techniques for MDP [camacho2019ltl, quint2021formal, li2018automata, li2019formal, tasse2020booleantaskalgebrareinforcement, jothimurugan2021compositional, hasanbeig2019logicallyconstrainedreinforcementlearning, hasanbeig2020cautious, hasanbeig2020deep, yalcinkaya2023automata, xu2024generalization, yalcinkaya2024compositional, yalcinkaya2025automataconditionedrl, jackermeier2025deepltl, shah2025ltlconstrained]. Within these works, there are two subcategories including methods for learning policies that can perform well across a set of multiple LTL tasks [vaezipoor2021ltl2action, yalcinkaya2023automata, yalcinkaya2024compositional, jackermeier2025deepltl] and methods that focus on optimally satisfying a single LTL task [toroicarte2018RewardMachines, camacho2019ltl, quint2021formal, li2019formal, hasanbeig2019logicallyconstrainedreinforcementlearning, hasanbeig2020cautious, hasanbeig2020deep, toroicarteRewardMachines2022, shah2025ltlconstrained].We specifically aim to improve on the scalability of these latter methods in high-dimensional environments. Of these techniques, we specifically highlight Reward Machines [toroicarteRewardMachines2022] as a framework with similar generality to our method and the first baseline we compare against.
@@
Our work utilizes techniques from GCRL [kaelbling1993learning, schaul2015universal] to accelerate learning for LTL tasks. Other methods have taken a related approach by learning or re-using goal-conditioned policies that are hierarchically guided by LTL expressions to enable zero-shot generalization to new tasks [qiu2023instructing, yalcinkaya2023automata, jackermeier2025deepltl, yalcinkaya2025automataconditionedrl]. Using discrete skills as opposed to a goal-conditioned policy has also been explored in [tasse2022skill, araki2021logicaloptionsframework]. All these approaches require the user to train a different skill for every automaton edge or a different goal-conditioned policy for every safety constraint that appears in the task, except for limited classes of safety constraints. This is impractical for most real-world tasks and, in particular, when safety constraints are changing throughout a task. Conversely, our method learns a single goal-conditioned policy for all goals in a particular LTL expression while still accounting for arbitrary, potentially non-stationary safety constraints. While ACQL has been developed for online RL rather than zero-shot generalization, there is a significant overlap in the tasks that both can ultimately solve. Therefore, we include the Logical Options Framework (LOF) [araki2021logicaloptionsframework] as a representative for these methods in our experiments.
@@
Our work is also focused on solving LTL tasks that feature safety constraints. Safety constraints are often indirectly addressed in the methods discussed above. They primarily focus on making unsafe actions suboptimal by either terminating rollouts [toroicarte2018RewardMachines, camacho2019ltl, toroicarteRewardMachines2022, hasanbeig2019logicallyconstrainedreinforcementlearning] or applying negative reward shaping [araki2021logicaloptionsframework, quint2021formal, vaezipoor2021ltl2action] when an unsafe automaton transition occurs. Other approaches based on formal methods, such as shielding [alshiekh2018shielding, konighofer2023online] and runtime model-checking [desai2017combining], can more robustly enforce safety requirements but typically rely on additional assumptions such as access to a discrete abstraction of the environment or known system dynamics. In contrast, techniques from the Safe RL literature can still robustly enforce stationary safety constraints in complex environments without such assumptions [altman1999constrained, yu2022reachabilityconstrainedreinforcementlearning, kalweit2020deepconstrainedqlearning, stooke2020responsivesafetyreinforcementlearning, ha2021saclagrangian, achiam2017cpo, liu2020ipo]. Our approach builds on these Safe RL methods to address non-stationary safety constraints induced by LTL specifications.
@@ 3 Preliminaries
RL problems with constraints are typically modeled as discounted CMDP, which are defined by a tuple $(\mathcal{S},\mathcal{A},\mathcal{T},d_{0},r,c,\mathcal{L},\gamma)$. This tuple consists of a state space $\mathcal{S}$, an action space $\mathcal{A}$, a transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow P(\mathcal{S})$ (where $P(\mathcal{S})$ represents the space of probability measures over $\mathcal{S}$), an initial state distribution $d_{0}\in P(\mathcal{S})$, reward and constraint functions $r,\;c:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, a discount factor $\gamma$, and a limit for constraint violation $\mathcal{L}\in\mathbb{R}$ [altman1999constrained]. Given a CMDP, the standard objective is to find the stationary policy $\pi:\mathcal{S}\to P(\mathcal{A})$ satisfying the constrained maximization
$\max_{\pi}\;J_{r}(\pi)\quad\textnormal{s.t.}\quad J_{c}(\pi)<\mathcal{L},$ (1)
@@
LTL [pnueli1977ltl] is an extension of propositional logic for reasoning about systems over time. LTL formulae $\phi\in\Phi$ consist of atomic propositions $p$ from a set $AP$, the standard boolean operators “not” ($\neg$), “and” ($\wedge$), and “or” ($\vee$), and temporal operators that reference the value of propositions in the future. These operators are “next” ($\circ$), “eventually” ($\lozenge$), and “always” ($\square$). We follow the definitions for these operators given in [baier2008principles]. In our algorithm, we use Signal Temporal Logic (STL), which further extends LTL with quantitative semantics and requires that every atomic proposition $p\in AP$ is defined as a real-valued function of the system state. The quantitative semantics is defined by a function $\rho:\mathcal{S}^{\omega}\times\Phi\rightarrow\mathbb{R}$ that produces a robustness value representing how much a sequence of states $\omega\in\mathcal{S}^{\omega}$ satisfies or violates the property specified by an STL formula $\phi\in\Phi$. Our exact implementation of these quantitative semantics is based on [dawson2022robust_stl_planning]. Every temporal logic constraint can be expressed as the conjunction of a “safety” and “liveness” constraint [alpern1985liveness], where a safety constraint is informally defined as a requirement that something must never happen for the task to succeed, and a liveness constraint is defined as a requirement that something must happen for the task to succeed. We take advantage of this dichotomy in our method.
@@
The robotics tasks discussed so far can be expressed more specifically as STL formulae in the recurrence class, as defined in [manna1990hierarchy], and all such formulae can be translated into abstract machines called Deterministic Büchi Automata [baier2008principles]. A DBA is formally defined by the tuple $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$, consisting of an alphabet $\Sigma$, a set of internal states $\mathcal{Q}$, a transition function $\delta:\mathcal{Q}\times\Sigma\rightarrow\mathcal{Q}$, an initial state $q_{0}\in\mathcal{Q}$, and a set of accepting states $F\subseteq\mathcal{Q}$. An automaton is used to process an infinite sequence of arbitrary symbols from the alphabet $\Sigma$ and determine whether the sequence satisfies or does not satisfy the original logical expression. In our setting, a symbol processed by the automaton is a subset of atomic propositions $l\in 2^{AP}$ that are true in some MDP state $s\in\mathcal{S}$. This symbol $l$ is referred to as the state’s labeling and the mapping of states to their labeling is denoted by a labeling function $L:\mathcal{S}\to 2^{AP}$. The automaton $A$ is always in some state $q\in\mathcal{Q}$, and starts in the state $q_{0}$. Each element $\sigma\in\Sigma$ of an input sequence causes some change in the automaton’s internal state according to the transition function $\delta$. For convenience, we will refer to edges in the automaton with transition predicates. A transition predicate is a propositional formula that holds for all $\sigma\in\Sigma$ which induce a transition between two states $q_{i},q_{j}\in\mathcal{Q}$ according to $\delta$ [alpern1987Recognizing]. By processing each state while an agent interacts with an MDP, the internal automaton state provides information to the agent about its current progress in the task. This is the fundamental idea behind a Product MDP, which combines any MDP with a deterministic automaton and is a standard technique in model checking for probabilistic systems [baier2008principles]. A complete definition for a product MDP in an RL context can be found in previous work such as [hasanbeig2020deep, toroicarteRewardMachines2022].
@@ 4 Method
Figure 1: The ACQL algorithm relies on a novel augmented formulation of CMDP (left). An input task specification $\phi$ is converted into the DBA AA, from which safety constraints and subgoals are collected into mappings $S$ and $G$ respectively. The learning agent receives subgoals $g_{1},\cdots,g_{n}$ at every stage in the task from $G$ and safety constraint feedback in $c^{A}$ from $S$. Trajectories induced by the policy $\pi_{j}$ are collected into a replay buffer $R$, from which batches $\mathcal{B}^{\tau}_{j}$ are sampled and modified using HER [andrychowicz2017her]. From these modified trajectories, mini-batches $\mathcal{B}_{i}$ of transitions $(s^{A}_{t},a_{t},r^{A}_{t},c^{A}_{t},s^{A}_{t+1})$ are used to compute the targets $y_{t}^{r}$ in (3) and $y_{t}^{c}$ in (4) for training models of the state-action value function and safety function, $Q^{r}_{\theta}$ and $Q^{c}_{\psi}$, from which an updated policy $\pi_{j+1}$ is derived.
@@ 4.1 Augmented Product CMDP Formulation
To construct the augmented product CMDP used by our algorithm, we begin by translating the input STL task specification $\phi$ into a DBA $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$ using the SPOT library [duret22cav]. As described in Section 3, this automaton represents the structure of the task. To better guide learning, we extract from this automaton additional structure that captures the task’s safety and liveness requirements. Intuitively, these represent conditions that must always hold to avoid failure (safety) and conditions that must eventually hold to make progress toward task completion (liveness) [alpern1985liveness, alpern1987Recognizing]. This information is summarized in two mappings: a safety condition mapping $S:\mathcal{Q}\rightarrow\Phi$, which assigns to each automaton state a proposition that must remain true, and a liveness condition mapping $O:\mathcal{Q}\rightarrow\Phi$, which assigns to each automaton state a proposition that must be eventually satisfied to proceed in the task. Based on $O$, we define $G:\mathcal{Q}\rightarrow\mathcal{G}^{+}$ to explicitly link automaton states to subgoals, where $\mathcal{G}^{+}=\bigcup_{i=1}^{n}\mathcal{G}^{i}$ represents the full space of possible subgoal lists up to a task-dependent length $n$ and $\mathcal{G}^{i}$ is the $i$-fold cartesian product of $\mathcal{G}$ with itself.
@@
Second, to obtain liveness constraints, we ignore the above transitions for safety conditions and examine the remaining predicates guarding outgoing transitions from each state $q$. These can only specify the conditions necessary to make progress in the task. Therefore, we can take the disjunction of these transition predicates to obtain the formula $O(q)$ that must be satisfied in order to transition beyond $q$. Following prior work on learning LTL tasks [araki2021logicaloptionsframework, qiu2023instructing], we assume that a subset of the atomic propositions, $AP_{\text{subgoal}}\subseteq AP$, represent subgoal propositions and are parameterized by values $g$ within a subgoal space $\mathcal{G}\subseteq\mathcal{S}$. Note that we consider only a subset since not all the propositions in our modified DBA are necessarily related to achieving subgoals. Using this fact, we filter the predicates in $O(q)$ to retain only those associated with these subgoal propositions, which in turn provides a list of subgoals $G(q)=g_{1},\dots,g_{n}$ that are relevant to the logical formula given by $O(q)$. An illustrative example for these two steps is provided in the appendix.
@@
Now, let $\mathcal{M}^{A}=(\mathcal{S}^{A},\mathcal{A}^{A},\mathcal{T}^{A},d_{0}^{A},r^{A},c^{A},\gamma,\mathcal{L})$ be a new CMDP formed by augmenting the original MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},d_{0},r,\gamma)$ with the DBA $A=(\Sigma,\mathcal{Q},\delta,q_{0},F)$. The new state space $\mathcal{S}^{A}=\mathcal{S}\times\mathcal{G}^{+}\times\mathcal{Q}$ is the Cartesian product of the original state space $\mathcal{S}$, the space for every possible list of subgoals $\mathcal{G}^{+}$, and the set of automaton states $\mathcal{Q}$. A state $s^{A}\in\mathcal{S}^{A}$ can be written as $\langle s,g^{+},q\rangle$, where $s\in\mathcal{S}$, $g^{+}\in\mathcal{G}^{+}$, and $q\in\mathcal{Q}$ are the constituent MDP state, goal-list and automaton state of $s^{A}$. The action space $\mathcal{A}^{A}=\mathcal{A}$ is unmodified. The new transition dynamics $\mathcal{T}^{A}$ are defined so that a transition to $\langle s^{\prime},g^{+}{{}^{\prime}},q^{\prime}\rangle\in\mathcal{S}^{A}$ from $\langle s,g^{+},q\rangle\in\mathcal{S}^{A}$ is impossible if the automaton does not support a transition from $q$ to $q^{\prime}$ when entering the state $s^{\prime}$; i.e.,
$\mathcal{T}^{A}$ $(\langle s^{\prime},g^{+}{{}^{\prime}},q^{\prime}\rangle|\langle s,g^{+},q\rangle,a)=\begin{cases}\mathcal{T}(s^{\prime}|s,a)&\text{if}\ q^{\prime}=\delta(q,L(s^{\prime})),g^{+}{{}^{\prime}}=G(q^{\prime}),\\ 0&\text{otherwise}.\end{cases}$